# Toward Better Temporal Structures for Geopolitical Events Forecasting

### v1

@@ 7 Conclusion
In this work, we presented HTKGH, an extension of HTKG that is designed to efficiently and accurately express complex facts involving many primary entities. Then, using the HTKGH’s formalization, we introduced the htkgh-polecast dataset, built on top of the POLECAT event database. Moreover, we conducted a thorough investigation into using state-of-the-art LLMs on a relation prediction task defined on htkgh-polecast. Our results showcased the out-of-the-box flexibility of LLMs to adapt to reasoning over complex higher-order facts, paving the way for future applications on even more complex structures and underlying data.
@@
In this work, we focused on LLMs due to their flexibility and ease of use out of the box, but it is possible to investigate various existing and novel architectures based on the GNNs’ framework. While Section 6 and Appendix K present a small study on a few GNN-based models, an in-depth and comprehensive experimentation with these models is warranted, which we leave to future work, as it is outside the scope of this work.
@@
One of the often overlooked criteria for graph structures is the knowledge representation perspective, which concerns itself with querying time and complexity using technologies such as SPARQL. Ideally, we want structures that both benefit learning algorithms and querying technology, demanding studies similar to the work by 10.1007/978-3-031-47240-4_15 for static knowledge graphs. Given the limited scope of this work, we leave such investigations to future work.
@@ C.3 Test Set Construction
Publisher Model #Params (B) Type
Google gemma-3-270m-it 0.3 N
gemma-3-1b-it 1.0 N
gemma-3-4b-it 4.0 N
gemma-3-12b-it 12.0 N
Qwen Qwen3-0.6B 0.8 H
Qwen3-1.7B 2.0 H
Qwen3-4B-Instruct-2507 4.0 N
Qwen3-4B-Thinking-2507 4.0 N
Qwen3-8B 8.0 H
Qwen3-14B 15.0 H
Meta Llama-3.1-8B-Instruct 8.0 N
DeepSeek DeepSeek-R1-Distill-Qwen-7B 8.0 T
OpenAI gpt-oss-20b 22.0 T
@@ F.2 Thinking
Appendix G Response Parser
@@
Appendix H Models
Table 3 showcases all 13 models used in our experiments.
@@ Appendix I Implementation Details
Appendix J Effect of Model Size
Figure 11 and Figure 12 present the full results of our experiments on the effect of model size on relation prediction accuracy for the Gemma-3 and Qwen3 families. In both families, we observe that performance generally increases as models get larger, with some minor exceptions in the smaller < 1B models. Notably, the exceptions happen in the scenarios with fewer contextual samples and looser filters, which showcases a potential lack of adequate information for reasoning, leading to a virtual upper bound in performance. Moreover, in the Qwen-3 family, we observe that the thinking variants outperform non-thinking ones consistently, with the 4B models (i.e., Qwen3-4B-*-2507) outperforming the rest of the family.
Appendix K Graph Neural Network Models
In this section, we study GNN-based models.
K.1 Preliminaries
@@
K.2 Fact Representation
For a given fact $\psi\in\Psi_{i}(q)$, we compute three role-specific DeepSet zaheer2017deep representations for actors, recipients, and qualifiers as
@@
where $f_{\psi}$ and $f_{r}$ are MLPs, and $\alpha_{\psi}$ and $\beta_{\psi}$ impose a relation-specific affine transformation (perez2018film). For query facts, where we do not have access to relation information, we use a placeholder “query" relation to avoid target leakage.
@@
where $f_{e}^{\tau}$ is a type-specific MLP. Finally, we use an aggregation scheme similar to R-GCN (schlichtkrull2018modeling) to update the node representation for $\psi\in\Psi_{i}(q)$ as
@@
where $f_{t}$ is a transformer (vaswani2017attention).
K.5 Linear Classifier
@@
K.6 Loss Functions
@@
K.8 Window Encoder Bottleneck Discussion
@@
K.9 Ablation on Number of Windows

### v2

@@ 7 Conclusion
In this work, we presented HTKGH, an extension of HTKG that is designed to efficiently and accurately express complex facts involving many primary entities. Then, using the HTKGH’s formalization, we introduced the htkgh-polecast dataset, built on top of the POLECAT event database. Moreover, we conducted a thorough investigation into using state-of-the-art LLMs on relation prediction and link prediction tasks defined on htkgh-polecast. Our results showcased 1) the out-of-the-box flexibility of LLMs to adapt to reasoning over complex higher-order facts and 2) the significant performance and inference time advantage HTKGH provides compared to existing structures, paving the way for future applications on even more complex structures and underlying data.
@@
In this work, we focused on LLMs due to their flexibility and ease of use out of the box, but it is possible to investigate advanced existing and novel architectures based on the GNNs’ framework, which might also solve the bottleneck issues we faced. While Section 6 and Appendix M present a small study on a few GNN-based models, an in-depth and comprehensive experimentation with these models is warranted, which we leave to future work.
Link Prediction
Given the aforementioned nuances regarding metrics and prediction strategies, there is a need for further investigations into the link prediction task. While we present a limited set of experiments for the impact of HTKGH on the link prediction task in Section 6, we leave further explorations to future work.
@@
One of the often overlooked criteria for graph structures is the knowledge representation perspective, which concerns itself with querying time and complexity using technologies such as SPARQL. Ideally, we want structures that both benefit learning algorithms and querying technology, demanding studies similar to the work by Iglesias-Molina et al. (2023) for static knowledge graphs. Given the limited scope of this work, we leave such investigations to future work.
@@ Appendix C Dataset
To create our dataset, rather than using raw text and parsing it ourselves, we utilize the already parsed data in the POLECAT database, which is stored in tabular format, and conform it to our HTKGH structure for our experiments.
@@ F.2 Thinking
Publisher Model #Params (B) Type
Google gemma-3-4b-it 4.0 N
gemma-3-12b-it 12.0 N
Qwen Qwen3-4B-Instruct-2507 4.0 N
Qwen3-4B-Thinking-2507 4.0 N
Qwen3-8B 8.0 H
Qwen3-14B 15.0 H
Meta Llama-3.1-8B-Instruct 8.0 N
DeepSeek DeepSeek-R1-Distill-Qwen-7B 8.0 T
OpenAI gpt-oss-20b 22.0 T
@@
Appendix G Models
Table 3 showcases all 9 models used in our experiments.
Appendix H Response Parser
@@ Appendix I Implementation Details
Appendix J Unidirectional vs. Bidirectional Edges
$(\{v\},r_{u},\{u\},t),Q)\;\;\text{vs.}\;\;(\{v,u\},r_{b},\{\},t),Q)$
Appendix K Effect of Model Size
While many training factors are often undisclosed, model size has always been one of the consistent predictors of performance. Figure 12 and Figure 13 showcases a comparison with regard to model size within the Gemma-3 and Qwen3 families. In both families, the performance generally increases as model sizes increase. The most significant exceptions are the Qwen3-4B-*-2507 models, which even beat the much larger Qwen3-14B model, underscoring the critical nature of other factors that must be studied for model selection.
Appendix L Effect of Number of Contextual Samples
One of the critical considerations of using LLMs for TKG-related tasks is the number of facts that could be included in the context. Figure 11 illustrates our experimental results with respect to the number of contextual samples. As is evident, most models generally benefit from the inclusion of more context, DeepSeek-R1-Distill-Qwen-7B being the only model that degrades with more samples. Moreover, we observe that scenarios with stricter filtering benefit more from an increase in samples. These results illustrate a promising trend in context scalings for more complex TKG-related predictions. Note that some models allow us to scale the samples further, but due to resource constraints, we leave these experiments to future work.
Appendix M Graph Neural Network Models
M.1 Preliminaries
@@
M.2 Fact Representation
For a given fact $\psi\in\Psi_{i}(q)$, we compute three role-specific DeepSet (Zaheer et al., 2017) representations for actors, recipients, and qualifiers as
@@
where $f_{\psi}$ and $f_{r}$ are MLPs, and $\alpha_{\psi}$ and $\beta_{\psi}$ impose a relation-specific affine transformation (Perez et al., 2018). For query facts, where we do not have access to relation information, we use a placeholder “query" relation to avoid target leakage.
@@
where $f_{e}^{\tau}$ is a type-specific MLP. Finally, we use an aggregation scheme similar to R-GCN (Schlichtkrull et al., 2018) to update the node representation for $\psi\in\Psi_{i}(q)$ as
@@
where $f_{t}$ is a transformer (Vaswani et al., 2017).
M.5 Linear Classifier
@@
Figure 15: Relation prediction accuracy (%) between the state-of-the-art open source and closed source thinking models.
M.6 Loss Functions
@@
M.8 Window Encoder Bottleneck Discussion
@@
M.9 Ablation on Number of Windows
@@
Appendix N Closed Source vs. Open Source Models
Figure 15 presents a comparison between the state-of-the-art open source and closed source thinking models. Surprisingly, the open source model (i.e., Qwen3-4B-Thinking-2507) edges out the closed source model (i.e., gpt-5.2) by +1.4%, showcasing an impressive performance on this task.